Q1) for a-firm how does the concept of producer surplus differ from that of profit.?

Q2) if the supply curve is q=3+1.5p what is the producer surplus if p=12?


1
Expert's answer
2021-10-31T18:28:46-0400

a.

Total revenues are subtracted from total costs to arrive at profit. Producer surplus, on the other hand, is the difference between the revenue from one item's sale and its marginal, direct cost of production - that is, the rise in overall cost produced by that item. Profit and producer surplus are the same when a business's sole costs are marginal, direct costs. However, if you have permanent or sunk expenditures, such as rent or new equipment, those costs are reflected into the profit calculation, and profit is less than producer surplus.

Assume you spend $10,000 on a frame-making machine. You're going to sell these photo frames for ten dollars apiece. Each photo frame costs $6 to manufacture, including supplies and direct labour. You get $4 in producer surplus for every sale. However, it would be incorrect to claim that your first sale resulted in a profit of $4. In actuality, you're still owing $9,996. You haven't recouped the cost of the machine until you sell your 2,500th picture frame. After then, every sale contributes to your profit.

b.

"Q=3+1.5p\\P=12\\Q=3+1.5(12)\\Q=21"

It will be a deficit of 75

Producer surplus

"=12\times 21-\int ^{21}_0[\frac{q^2-2q}{}-2q]^{21}_{0}\\=252-(147-42)\\=147"

producer surplus is 147 units